OpenStudy (anonymous):

Product of powers property X4.x3=___x(12)___

jimthompson5910 (jim_thompson5910):

Rule: \[\LARGE x^a*x^b = x^{a+b}\]

OpenStudy (anonymous):

I don't get is that all

OpenStudy (unklerhaukus):

\[2^3=\overbrace{2\times 2\times2}^{3\text{ times}}\] \[x^a = \overbrace{x\times x\times\cdots\times x}^{a\text{ times}}\] \[x^a\times x^b = \overbrace{x\times x\times\cdots\times x}^{a\text{ times}}\times \overbrace{x\times x\times\cdots\times x}^{b\text{ times}}\\ \qquad\quad =\overbrace{x\times x\times\cdots\times x}^{a+b\text{ times}}\\ \qquad\quad =x^{a+b}\]

OpenStudy (unklerhaukus):

\[(x^a)^b = \overbrace{\overbrace{x\times x\times\cdots\times x}^{a\text{ times}}\times\overbrace{x\times x\times\cdots\times x}^{a\text{ times}}\times\cdots \overbrace{x\times x\times\cdots\times x}^{a\text{ times}}}^{b\text{ times}}\\ \quad\quad= x^{a\times b}\]

OpenStudy (anonymous):

So what's the answer

OpenStudy (unklerhaukus):

Can you write out your question again please? (it's not exactly clear)

OpenStudy (anonymous):

X4•X3=__x(12)__

OpenStudy (unklerhaukus):

\[x^4\cdot x^3 = ?x^{12}?\]

OpenStudy (anonymous):

Yes

OpenStudy (unklerhaukus):

what are the underscores/blanks??

OpenStudy (unklerhaukus):